# Learning multi-modal generative models with permutation-invariant encoders and tighter variational bounds

### v2

@@ 2.2 Information-theoretic perspective
Beyond generative modeling, $\beta$-VAEs (Higgins et al., 2017) have been popular for representation learning and data reconstruction. Alemi et al. (2018) suggest learning a latent representation that achieves certain mutual information with the data based on upper and lower variational bounds of the mutual information. A Legendre transformation thereof recovers the $\beta$-VAE objective and allows a trade-off between information content or rate versus reconstruction quality or distortion. We show that the proposed variational objective gives rise to an analogous perspective for multiple modalities. Recall that the mutual information on the inference path [^4] We include the conditioning modalities as an index for the latent variable $Z$ when the conditioning set is unclear. is given by
@@
can be bounded by standard (Barber and Agakov, 2004; Alemi et al., 2016, 2018) lower and upper bounds:
$\mathcal{H}_{\mathcal{S}}-D_{\mathcal{S}}\leq\mathcal{H}_{\mathcal{S}}-D_{\mathcal{S}}+\Delta_{1}=\operatorname{I}_{q_{\phi}}(X_{\mathcal{S}},Z)=R_{\mathcal{S}}-\Delta_{2}\leq R_{\mathcal{S}},$ (6)
with $\Delta_{1},\Delta_{2}\geq 0$ for the rate
measuring the information content that is encoded by $q_{\phi}$ into the latents, and the distortion
$D_{\mathcal{S}}=-\int q_{\phi}(x_{\mathcal{S}},z)\log p_{\theta}(x_{\mathcal{S}}|z)\mathrm{d}z\mathrm{d}x_{\mathcal{S}}$
given as the negative reconstruction log-likelihood. For details, see Appendix C. Observe that for any $\beta>0$ by construction,
$-\int p_{d}(x_{\mathcal{S}})\mathcal{L}(x_{\mathcal{S}})\mathrm{d}x_{\mathcal{S}}=D_{\mathcal{S}}+\beta R_{\mathcal{S}}$
and for any and that $\mathcal{H}_{\mathcal{S}}\leq R_{\mathcal{S}}+D_{\mathcal{S}}$. To arrive at a similar interpretation for the conditional bound $\mathcal{L}_{\setminus\mathcal{S}}$ that involves the conditional mutual information
@@
Using the chain rules for entropy, we obtain that the suggested bound can be seen as a relaxation of bounds on marginal and conditional mutual information.
@@
and minimizing $\mathcal{L}$ for fixed $\beta=\frac{\partial(D_{\mathcal{S}}+D_{\setminus\mathcal{S}})}{\partial(R_{\mathcal{S}}+R_{\setminus\mathcal{S}})}$ minimizes the rates $R_{\mathcal{S}}+R_{\setminus\mathcal{S}}$ and distortions $D_{\mathcal{S}}+D_{\setminus\mathcal{S}}$.
Remark 9 (Mixture based variational bound)
The arguments in Daunhawer et al. (2022) imply that
$-\int p_{d}(\mathrm{d}x)\mathcal{L}^{\text{Mix}}_{\mathcal{S}}(x)=D_{\mathcal{S}}+D^{c}_{\setminus\mathcal{S}}+\beta R_{\mathcal{S}},$
where
$D^{c}_{\setminus\mathcal{S}}=-\int p_{d}(x_{\mathcal{S}})q_{\phi}(z_{\mathcal{S}}|x_{\mathcal{S}})\log p_{\theta}(x_{\setminus\mathcal{S}}|z_{\mathcal{S}})\mathrm{d}z_{\mathcal{S}}\mathrm{d}x_{\mathcal{S}}$
is a cross-distortion term. Due to
$\mathcal{H}(X_{\mathcal{M}}|Z_{\mathcal{S}})=-\mathcal{H}(X_{\mathcal{M}})+\operatorname{I}_{q_{\phi}}(X_{\mathcal{M}},Z_{\mathcal{S}})\leq D_{\mathcal{S}}+D^{c}_{\setminus\mathcal{S}}$
we can view minimizing $\mathcal{L}^{\text{Mix}}_{\mathcal{S}}$ as minimizing
$\mathcal{H}(X_{\mathcal{M}})-\operatorname{I}_{q_{\phi}}(X_{\mathcal{M}},Z_{\mathcal{S}})+\beta\operatorname{I}_{q_{\phi}}(X_{\mathcal{S}},Z_{\mathcal{S}}),$
see (2).
@@
see also Huang et al. (2020) and Remark 21. Similarly, if (7) holds, then it is readily seen that the minimum of the bound $\int p_{d}(\mathrm{d}x)\mathcal{L}_{\setminus\mathcal{S}}(x)$ is attained at any $x$ for the variational density $q^{\star}(z|x)=\tilde{p}_{\beta,\theta}(z|x)$. In contrast, as shown in Appendix 22, the optimal variational density for the mixture-based (1) multi-modal objective is attained at
@@ 3 Permutation-invariant modality encoding
  • Mixture of Product of Experts (MoPoE), see Sutter et al. (2021),
    $q_{\varphi}^{\text{MoPoE}}(z|x_{\mathcal{S}})=\frac{1}{|\mathcal{P}(\mathcal{S})|)}\sum_{\mathcal{A}\subset\mathcal{S}}q_{\varphi}^{\text{MoE}}(z|x_{\mathcal{A}})$
@@ 3.1 Learnable permutation-invariant aggregation schemes
We aim to learn a more flexible aggregation scheme under the constraint that the encoding distribution is invariant (Bloem-Reddy and Teh, 2020) with respect to the ordering of encoded features of each modality. Put differently, for all $(H_{s})_{s\in\mathcal{S}}\in\mathbb{R}^{|\mathcal{S}|\times D_{E}}$ and all permutations $\pi\in\mathbb{S}_{\mathcal{S}}$ of $\mathcal{S}$, we assume that the conditional distribution is $\mathbb{S}_{\mathcal{S}}$-invariant, i.e. $q^{\prime}_{\vartheta}(z|h)=q^{\prime}_{\vartheta}(z|\pi\cdot h)$ for all $z\in\mathbb{R}^{D}$, where $\pi$ acts on $H=(H_{s})_{s\in\mathcal{S}}$ via $\pi\cdot H=(H_{\pi(s)})_{s\in\mathcal{S}}$. We set $q_{\phi}(z|x_{\mathcal{S}})=q^{\prime}_{\vartheta}(z|{h_{s,\varphi}(x_{s})}_{s\in\mathcal{S}})$, $\phi=(\varphi,\vartheta)$ and remark that the encoding distribution is not invariant with respect to the modalities, but becomes only invariant after applying modality-specific encoder functions $h_{s,\varphi}$. Observe that such a constraint is satisfied by the aggregation schemes above for $h_{s,\varphi}$ being the uni-modal encoders.
A variety of invariant (or equivariant) functions along with their approximation properties have been considered previously, see for instance Santoro et al. (2017); Zaheer et al. (2017); Qi et al. (2017); Lee et al. (2019); Segol and Lipman (2019); Murphy et al. (2019); Maron et al. (2019); Sannai et al. (2019); Yun et al. (2019); Bruno et al. (2021); Wagstaff et al. (2022); Zhang et al. (2022b); Li et al. (2022); Bartunov et al. (2022), and applied in different contexts such as meta-learning (Edwards and Storkey, 2016; Garnelo et al., 2018b; Kim et al., 2018; Hewitt et al., 2018; Giannone and Winther, 2022), reinforcement learning (Tang and Ha, 2021; Zhang et al., 2022a) or generative modeling of (uni-modal) sets (Li et al., 2018, 2020; Kim et al., 2021; Biloš and Günnemann, 2021; Li and Oliva, 2021). We can use such constructions to parameterize more flexible encoding distributions. Indeed, the results from Bloem-Reddy and Teh (2020) imply that for an exchangable sequence $H_{\mathcal{S}}=(H_{s})_{s\in\mathcal{S}}\in\mathbb{R}^{|\mathcal{S}|\times D_{E}}$ and random variable $Z$, the distribution $q^{\prime}(z|h_{\mathcal{S}})$ is $\mathbb{S}_{\mathcal{S}}$-invariant if and only if there is a measurable function [^5] The function $f^{\star}$ generally depends on the cardinality of $\mathcal{S}$. Finite-length exchangeable sequences imply a de Finetti latent variable representation only up to approximation errors (Diaconis and Freedman, 1980). $f^{\star}\colon[0,1]\times\mathcal{M}(\mathbb{R}^{D_{E}})\to\mathbb{R}^{D}$ such that
@@
The Deep Set (Zaheer et al., 2017) construction $f_{\vartheta}(h_{\mathcal{S}})=\rho_{\vartheta}\left(\sum_{s\in\mathcal{S}}\chi_{\vartheta}(h_{s})\right)$ applies the same neural network $\chi_{\vartheta}\colon\mathbb{R}^{D_{E}}\to\mathbb{R}^{D_{P}}$ to each encoded feature $h_{s}$. We assume that $\chi_{\vartheta}$ is a feed-forward neural network and remark that pre-activation ResNets (He et al., 2016) have been advocated for deeper $\chi_{\vartheta}$. For exponential family models, the optimal natural parameters of the posterior solve an optimization problem where the dependence on the generative parameters from the different modalities decomposes as a sum, see Appendix F.
@@
Let $\text{MTB}_{\vartheta}$ be a multi-head pre-layer-norm transformer block (Wang et al., 2019a; Xiong et al., 2020), see Appendix D for precise definitions. For some neural network $\chi_{\vartheta}\colon\mathbb{R}^{D_{E}}\to\mathbb{R}^{D_{P}}$, set $g^{0}_{\mathcal{S}}=\chi_{\vartheta}(h_{\mathcal{S}})$ and for $k\in\{1,\ldots,L\}$, set $g_{\mathcal{S}}^{k}=\text{MTB}_{\vartheta}(g_{\mathcal{S}}^{k-1})$. We then consider $f_{\vartheta}(h_{\mathcal{S}})=\rho_{\vartheta}\left(\sum_{s\in\mathcal{S}}g^{L}_{s}\right)$. This can be seen as a Set Transformer (Lee et al., 2019; Zhang et al., 2022a) model without any inducing points as for most applications, a computational complexity that scales quadratically in the number of modalities can be acceptable. In our experiments, we use layer normalization (Ba et al., 2016) within the transformer model, although, for example, set normalization (Zhang et al., 2022a) could be used alternatively.
Remark 11 (Context-aware pooling)
$o_{s}=\frac{1}{Z}\sum_{t\in\mathcal{S}}\kappa(g_{s},g_{t})v(g_{t}),$
can be seen as a learnable non-symmetric kernel (Wright and Gonzalez, 2021; Cao, 2021). Conceptually, the attention encoder pools a learnable $D$-dimensional function $v$ using a learnable context-dependent weighting function. While such attention models directly account for the interaction between the different encodings, a DeepSet aggregation approach may require a sufficiently high-dimensional latent space $D_{P}$ to achieve universal approximation properties (Wagstaff et al., 2022).
@@
Sutter et al. (2021) introduced a MoPoE aggregation scheme that extends MoE or PoE schemes by considering a mixture distribution of all $2^{M}$ modality subsets, where each mixture component consists of a PoE model, i.e.,
@@
Combining expert distributions has a long tradition in decision theory and Bayesian inference; see Genest and Zidek (1986) for early works, with popular schemes being linear pooling (i.e., MoE) or log-linear pooling (i.e., PoE with tempered densities). These are optimal schemes for minimizing different objectives, namely a weighted (forward or reverse) KL-divergence between the pooled distribution and the inidividual experts (Abbas, 2009). Log-linear pooling operators are externally Bayesian, allowing for consistent Bayesian belief updates when each expert updates her belief with the same likelihood function (Genest et al., 1986).
@@ 3.2 Permutation-equivariance and private latent variables
However, the specific generative model also lends itself to an alternative parameterization that guarantees that cross-modal reconstruction likelihoods from $X_{\setminus\mathcal{S}}$ do not affect the encoding distribution of $\tilde{Z}_{\mathcal{S}}$ under our new variational bound. The assumption of private latent variables suggests an additional permutation-equivariance into the encoding distribution that approximates the posterior in (10), in the sense that for any permutation $\pi\in\mathbb{S}_{\mathcal{S}}$, it holds that
@@
The results from Bloem-Reddy and Teh (2020) then imply, for fixed $|\mathcal{S}|$, the existence of a function $f^{\star}$ such that for all $s\in\mathcal{S}$, almost surely,
@@
Similar to previous work Wang et al. (2016); Lee and Pavlovic (2021); Sutter et al. (2020), we consider an encoding density of the form
@@
is a (permutation-invariant) PoE aggregation, and we assumed that the prior density factorizes over the shared and different private variables. For each modality $s$, we encode different features $h^{\prime}_{s,\varphi}=(\mu^{\prime}_{s,\varphi},\Sigma^{\prime}_{s,\varphi})$ and $\tilde{h}_{s,\varphi}=(\tilde{\mu}_{s,\varphi},\tilde{\Sigma}_{s,\varphi})$ for the shared, respectively, private, latent variables.
@@
Similar to a Sum-Pooling approach, we consider an encoding density that writes as
@@
Remark 14 (Cross-modal context variables)
In contrast to the PoE model, where the private encodings are independent, the private encodings are dependent in the Sum-Pooling model by conditioning on a sample from the shared latent space. The shared latent variable $Z^{\prime}$ can be seen as a shared cross-modal context variable, and similar probabilistic constructions to encode such context variables via permutation-invariant models have been suggested in few-shot learning algorithms (Edwards and Storkey, 2016; Giannone and Winther, 2022) or, particularly, for neural process models (Garnelo et al., 2018b, a; Kim et al., 2018). Permutation-equivariant models have been studied for stochastic processes where invariant priors correspond to equivariant posteriors (Holderrieth et al., 2021), such as Gaussian processes or Neural processes with private latent variables, wherein dependencies in the private latent variables can be constructed hierarchically (Wang and Van Hoof, 2020; Xu et al., 2023).

### v3

@@ 2.2 Information-theoretic perspective
Beyond generative modeling, $\beta$-VAEs [Higgins et al., 2017] have been popular for representation learning and data reconstruction. Alemi et al. [2018] suggest learning a latent representation that achieves certain mutual information with the data based on upper and lower variational bounds of the mutual information. A Legendre transformation thereof recovers the $\beta$-VAE objective and allows a trade-off between information content or rate versus reconstruction quality or distortion. We show that the proposed variational objective gives rise to an analogous perspective for multiple modalities. Recall that the mutual information on the inference path [^3] We include the conditioning modalities as an index for the latent variable $Z$ when the conditioning set is unclear. is given by
@@
can be bounded by standard [Barber and Agakov, 2004, Alemi et al., 2016, 2018] lower and upper bounds:
$\mathcal{H}_{\mathcal{S}}-D_{\mathcal{S}}\leq\mathcal{H}_{\mathcal{S}}-D_{\mathcal{S}}+\Delta_{1}=\operatorname{I}_{q_{\phi}}(X_{\mathcal{S}},Z_{\mathcal{S}})=R_{\mathcal{S}}-\Delta_{2}\leq R_{\mathcal{S}},$ (7)
with $\Delta_{1},\Delta_{2}\geq 0$ and $\mathcal{H}_{\mathcal{S}}\leq R_{\mathcal{S}}+D_{\mathcal{S}}$. For details, see Appendix C. Consequently, by tuning $\beta$, we can vary upper and lower bounds of $\operatorname{I}_{q_{\phi}}(X_{\mathcal{S}},Z_{\mathcal{S}})$ to tradeoff between compressing and reconstructing $X_{\mathcal{S}}$.
To arrive at a similar interpretation for the conditional bound $\mathcal{L}_{\setminus\mathcal{S}}$ that involves the conditional mutual information
@@
Consequently, by tuning $\beta$, we can vary upper and lower bounds of $\operatorname{I}_{q_{\phi}}(X_{\setminus\mathcal{S}},Z_{\mathcal{M}}|X_{\mathcal{S}})$ to tradeoff between compressing relative to ${q_{\phi}}(\cdot|x_{\mathcal{S}})$ and reconstructing $X_{\setminus\mathcal{S}}$. Using the chain rules for entropy, we obtain that the suggested bound can be seen as a relaxation of bounds on marginal and conditional mutual information.
@@
and maximizing $\mathcal{L}$ for fixed $\beta=\frac{\partial(D_{\mathcal{S}}+D_{\setminus\mathcal{S}})}{\partial(R_{\mathcal{S}}+R_{\setminus\mathcal{S}})}$ minimizes the rates $R_{\mathcal{S}}+R_{\setminus\mathcal{S}}$ and distortions $D_{\mathcal{S}}+D_{\setminus\mathcal{S}}$.
Remark 9 (Mixture-based variational bound)
We show in Appendix C, see also Daunhawer et al. [2022], that
$\mathcal{H}_{\mathcal{M}}-D_{\mathcal{S}}-D_{\mathcal{S}}^{c}\leq\mathcal{H}_{\mathcal{M}}-D_{\mathcal{S}}-D_{\mathcal{S}}^{c}+\Delta_{1}^{\prime}=\operatorname{I}_{q_{\phi}}(X_{\mathcal{M}},Z_{\mathcal{S}}),$
where $\Delta_{1}^{\prime}=\int q_{\phi}^{\text{agg}}(z)\mathsf{K}\mathsf{L}(q^{\star}(x|z)|p_{\theta}(x|z))\mathrm{d}z>0$. Consequently, $\mathcal{H}_{\mathcal{M}}-D_{\mathcal{S}}-D_{\mathcal{S}}^{c}$ is a variational lower bound, while $R_{\mathcal{S}}$ is a variational upper bound on $\operatorname{I}_{q_{\phi}}(X_{\mathcal{M}},Z_{\mathcal{S}})$, which establishes (2). Maximizing the mixture-based bound thus corresponds to encoding a single latent variable $Z_{\mathcal{S}}$ that maximizes the reconstruction of all modalities while at the same time being maximally compressive relative to the prior.
@@
see also Huang et al. [2020] and Remark 18. Similarly, if (8) holds, then it is readily seen that the minimum of the bound $\int p_{d}(\mathrm{d}x)\mathcal{L}_{\setminus\mathcal{S}}(x)$ is attained at any $x$ for the variational density $q^{\star}(z|x)=\tilde{p}_{\beta,\theta}(z|x)$. In contrast, as shown in Appendix 19, the optimal variational density for the mixture-based (1) multi-modal bound is attained at
@@
The optimal variational density for the mixture-based bound thus tilts the posterior distribution to points that achieve higher cross-modal predictions.
@@ 3 Permutation-invariant modality encoding
  • Mixture of Product of Experts (MoPoE), see Sutter et al. [2021],
@@ 3.1 Learnable permutation-invariant aggregation schemes
We aim to learn a more flexible aggregation scheme under the constraint that the encoding distribution is invariant [Bloem-Reddy and Teh, 2020] with respect to the ordering of encoded features of each modality. Put differently, for all $(H_{s})_{s\in\mathcal{S}}\in\mathbb{R}^{|\mathcal{S}|\times D_{E}}$ and all permutations $\pi\in\mathbb{S}_{\mathcal{S}}$ of $\mathcal{S}$, we assume that the conditional distribution is $\mathbb{S}_{\mathcal{S}}$-invariant, i.e. $q^{\prime}_{\vartheta}(z|h)=q^{\prime}_{\vartheta}(z|\pi\cdot h)$ for all $z\in\mathbb{R}^{D}$, where $\pi$ acts on $H=(H_{s})_{s\in\mathcal{S}}$ via $\pi\cdot H=(H_{\pi(s)})_{s\in\mathcal{S}}$. We set $q_{\phi}(z|x_{\mathcal{S}})=q^{\prime}_{\vartheta}(z|{h_{s,\varphi}(x_{s})}_{s\in\mathcal{S}})$, $\phi=(\varphi,\vartheta)$ and remark that the encoding distribution is not invariant with respect to the modalities, but becomes only invariant after applying modality-specific encoder functions $h_{s,\varphi}$. Observe that such a constraint is satisfied by the aggregation schemes above for $h_{s,\varphi}$ being the uni-modal encoders.
A variety of invariant (or equivariant) functions along with their approximation properties have been considered previously, see for instance Santoro et al. [2017], Zaheer et al. [2017], Qi et al. [2017], Lee et al. [2019], Segol and Lipman [2019], Murphy et al. [2019], Maron et al. [2019], Sannai et al. [2019], Yun et al. [2019], Bruno et al. [2021], Wagstaff et al. [2022], Zhang et al. [2022b], Li et al. [2022], Bartunov et al. [2022], and applied in different contexts such as meta-learning [Edwards and Storkey, 2016, Garnelo et al., 2018b, Kim et al., 2018, Hewitt et al., 2018, Giannone and Winther, 2022], reinforcement learning [Tang and Ha, 2021, Zhang et al., 2022a] or generative modeling of (uni-modal) sets [Li et al., 2018, 2020, Kim et al., 2021, Biloš and Günnemann, 2021, Li and Oliva, 2021]. We can use such constructions to parameterize more flexible encoding distributions. Indeed, the results from Bloem-Reddy and Teh [2020] imply that for an exchangable sequence $H_{\mathcal{S}}=(H_{s})_{s\in\mathcal{S}}\in\mathbb{R}^{|\mathcal{S}|\times D_{E}}$ and random variable $Z$, the distribution $q^{\prime}(z|h_{\mathcal{S}})$ is $\mathbb{S}_{\mathcal{S}}$-invariant if and only if there is a measurable function [^4] The function $f^{\star}$ generally depends on the cardinality of $\mathcal{S}$. Finite-length exchangeable sequences imply a de Finetti latent variable representation only up to approximation errors [Diaconis and Freedman, 1980]. $f^{\star}\colon[0,1]\times\mathcal{M}(\mathbb{R}^{D_{E}})\to\mathbb{R}^{D}$ such that
@@
The Deep Set [Zaheer et al., 2017] construction $f_{\vartheta}(h_{\mathcal{S}})=\rho_{\vartheta}\left(\sum_{s\in\mathcal{S}}\chi_{\vartheta}(h_{s})\right)$ applies the same neural network $\chi_{\vartheta}\colon\mathbb{R}^{D_{E}}\to\mathbb{R}^{D_{P}}$ to each encoded feature $h_{s}$. We assume that $\chi_{\vartheta}$ is a feed-forward neural network and remark that pre-activation ResNets [He et al., 2016] have been advocated for deeper $\chi_{\vartheta}$. For exponential family models, the optimal natural parameters of the posterior solve an optimization problem where the dependence on the generative parameters from the different modalities decomposes as a sum, see Appendix F.
@@
Let $\text{MTB}_{\vartheta}$ be a multi-head pre-layer-norm transformer block [Wang et al., 2019a, Xiong et al., 2020], see Appendix D for precise definitions. For some neural network $\chi_{\vartheta}\colon\mathbb{R}^{D_{E}}\to\mathbb{R}^{D_{P}}$, set $g^{0}_{\mathcal{S}}=\chi_{\vartheta}(h_{\mathcal{S}})$ and for $k\in\{1,\ldots,L\}$, set $g_{\mathcal{S}}^{k}=\text{MTB}_{\vartheta}(g_{\mathcal{S}}^{k-1})$. We then consider $f_{\vartheta}(h_{\mathcal{S}})=\rho_{\vartheta}\left(\sum_{s\in\mathcal{S}}g^{L}_{s}\right)$. This can be seen as a Set Transformer [Lee et al., 2019, Zhang et al., 2022a] model without any inducing points as for most applications, a computational complexity that scales quadratically in the number of modalities can be acceptable. In our experiments, we use layer normalization [Ba et al., 2016] within the transformer model, although, for example, set normalization [Zhang et al., 2022a] could be used alternatively.
Note that the PoE’s aggregation mechanism involves taking inverses, which can only be approximated by the learned aggregation models. The considered permutation-invariant models can thus only recover a PoE scheme under universal approximation assumptions.
@@
Sutter et al. [2021] introduced a MoPoE aggregation scheme that extends MoE or PoE schemes by considering a mixture distribution of all $2^{M}$ modality subsets, where each mixture component consists of a PoE model, i.e.,
@@
Combining expert distributions has a long tradition in decision theory and Bayesian inference; see Genest and Zidek [1986] for early works, with popular schemes being linear pooling (i.e., MoE) or log-linear pooling (i.e., PoE with tempered densities). These are optimal schemes for minimizing different objectives, namely a weighted (forward or reverse) KL-divergence between the pooled distribution and the individual experts [Abbas, 2009]. Log-linear pooling operators are externally Bayesian, allowing for consistent Bayesian belief updates when each expert updates her belief with the same likelihood function [Genest et al., 1986].
@@ 3.2 Permutation-equivariance and private latent variables
However, the specific generative model also lends itself to an alternative parameterization that guarantees that cross-modal likelihoods from $X_{\setminus\mathcal{S}}$ do not affect the encoding distribution of $\tilde{Z}_{\mathcal{S}}$ under our new variational objective. The assumption of private latent variables suggests an additional permutation-equivariance into the encoding distribution that approximates the posterior in (11), in the sense that for any permutation $\pi\in\mathbb{S}_{\mathcal{S}}$, it holds that
@@
The results from Bloem-Reddy and Teh [2020] then imply, for fixed $|\mathcal{S}|$, the existence of a function $f^{\star}$ such that for all $s\in\mathcal{S}$, almost surely,
@@
$q_{\phi}(z^{\prime},\tilde{z}_{\mathcal{M}}|x_{\mathcal{S}})=q_{\phi}(z^{\prime}|x_{\mathcal{S}})p_{\theta}(\tilde{z}_{\setminus\mathcal{S}}|z^{\prime})q_{\phi}(\tilde{z}_{\mathcal{S}}|z^{\prime},x_{\mathcal{S}})$
so that the modality-specific information of $x_{\mathcal{S}}$ as encoded via $\tilde{z}_{\mathcal{S}}$ is not impacted by the realisation $\tilde{Z}_{\setminus\mathcal{S}}$ of modality-specific variation from the other modalities.
@@
Similar to previous work Wang et al. [2016], Lee and Pavlovic [2021], Sutter et al. [2020], we consider an encoding density of the form
@@
is a (permutation-invariant) PoE aggregation, and we assumed that the prior density factorizes over the shared and different private variables. For each modality $s$, we encode different features $h^{\prime}_{s,\varphi}=(\mu^{\prime}_{s,\varphi},\Sigma^{\prime}_{s,\varphi})$ and $\tilde{h}_{s,\varphi}=(\tilde{\mu}_{s,\varphi},\tilde{\Sigma}_{s,\varphi})$ for the shared, respectively, private, latent variables. We followed previous works [Tsai et al., 2019b, Lee and Pavlovic, 2021, Sutter et al., 2020] in that the encodings and prior distributions for the modality-specific latent variables are independent of the shared latent variables. However, this assumption can be relaxed, as long as the distributions remain Gaussian.
@@
Similar to a Sum-Pooling approach, we consider an encoding density that is written as